OpenStudy (anonymous):

simplify (3x^2-4x+1) (2x+x^2-9)

OpenStudy (anonymous):

dude take help from these 1. simplify the expression x^2+4x+3/ x^2-9 A. x-3/x-3 x not equal to 3 or -3 B. x-1/x-3 x not equal to 3 or -3 C. x+1/x-3 x not equal to 3 or -3 D. none of these 2. Solve the equation 3^3x = 81^3x-4 A. 4/3 B. no solution C. 16/9 D. 4/3 3. 3^2x=9^2x+4 x=___? A. x = 4 B. x = 8 C. x = -4 D. x = 2 4. The vertical asymptote in the graph of the equation f(x)=7/x-3 would be the line: A. x=3 B. x=-3 C. y=3 D. y=-3 2 years ago Report Abuse 1. (x² + 4x + 3) / (x² - 9) = (x + 3)(x + 1) / (x - 3)(x + 3) = (x + 1) / (x - 3) Look back at the original equation, you have (x² - 9) in the denominator, this means that x cannot be 3 or -3. Answer is c). 2. 3^(3x) = 81^(3x - 4) 3^(3x) = [3^4]^(3x - 4) 3^(3x) = 3^(12x - 16) 3x = 12x - 16 9x = 16 x = 16 / 9 Answer is c). 3. 3^(2x) = 9^(2x + 4) 3^(2x) = [3²]^(2x + 4) 3^(2x) = 3^(4x + 8) 2x = 4x + 8 2x = -8 x = -4 Answer is c). 4. f(x) = 7 / (x - 3); set the denominator to 0 and solve for x to get vertical asymptote x - 3 = 0 x = 3

OpenStudy (anonymous):

You got to simplify both separately

OpenStudy (campbell_st):

Use a grid  add along the diagonals to collect like terms the answer is \[3x^4 + 2x^3 -34x^2 + 38x - 9\]